OpenStudy (anonymous):

radical equations help how do you solve √(x + 8) = √(3x+8)

OpenStudy (anonymous):

Square both sides

OpenStudy (anonymous):

$$x+8=3x+8 \Rightarrow x=0$$

OpenStudy (anonymous):

x+8 =3x +8 x-3x=8-8 -2x=0 x=0

OpenStudy (anonymous):

is it like a regular algebra problem where you're solving for x after you get rid of the radical?

OpenStudy (anonymous):

people